# Massively Parallel Minimum Spanning Tree in General Metric Spaces

### v1

@@ 5 A Lower Bound for Metric MST
Claim 5.3 follows easily from Claim 5.2.
Proof of Claim 5.3.
We are given a graph $G$, comprising $\frac{1}{\varepsilon^{k}}$ cycles $C_{0},\ldots,C_{1/\varepsilon^{k}-1}$ of length $\varepsilon^{k}n$, and one vertex $u^{i}_{0}$ per cycle. Using Claim 5.2, we compute for each vertex $u$, its $\frac{1}{\varepsilon}$-neighborhood. This neighborhood can be collected in one machine, i.e. the subpath of length $\frac{2}{\varepsilon}$ centered at the vertex is retrieved.
To create $G^{\prime}$, on the same vertex set, we connect each vertex to the two vertices at distance $\frac{1}{\varepsilon}$ from it in $G$. Since the length of the cycles is divisible by $\frac{1}{\varepsilon}$, each cycle $C_{i}$ is broken down into $\frac{1}{\varepsilon}$ cycles $C^{i}_{0},\ldots,C^{i}_{1/\varepsilon-1}$ of length $\varepsilon^{k+1}n$. Also, $u^{i}_{j}$ is already computed for $10\leq i\leq 1/\varepsilon^{k}-$ 1 and $10\leq j\leq 1/\varepsilon-$ 1 using Claim 5.2. It is the subpath starting at $u^{i}_{0}$. ∎
[…]
Appendix A Lower Bounds (LOCAL and Component Stable MPC)
In this section, we present lower bounds for MST approximation on $(1,2)$-metrics. First, we prove a lower bound for a slightly easier problem in the LOCAL model with shared randomness. Then, we use the framework of [GhaffariKU19] and [CzumajDP21a] to lift it and get a $\Omega\left(\log\frac{1}{\varepsilon}\right)$-round lower bound for component stable MPC algorithms that compute a $(1+\varepsilon)$-approximate MST in $(1,2)$-metrics.
We note that while the lower bound proved in this section is weaker than the lower bound of Theorem 2 which does not require component stability assumption, we state it nonetheless to contrast lower bounds obtained from this framework with our lower bound of Theorem 2.
@@ A.1 LOCAL Model
We restate some of the claims and definitions from Section 5. Given a graph $G$, the $(1,2)$-metric obtained from $G$ is such that for any two vertices $u$ and $v$, their distance in the metric is $1$ if and only if there is an edge between them in $G$. We say a spanning tree $T$ of the metric removes an edge $e\in G$ if $e\notin T$ (Definition 5.6). That is, $T$ is constructed by removing some edges from $G$, turning it into a forest, and then connecting the resulting components with weight-$2$ edges. When $G$ is a disjoint union of cycles, a $(1+\varepsilon)$-approximate MST of this metric removes at least one edge from each cycle, and at most $c+\varepsilon(n+c-2)$ edges in total, where $n$ is the number of vertices, and $c$ is the number of cycles (Claim 5.7).
@@
In the $\varepsilon$-almost-spanning forest problem, we are given a parameter $10<\varepsilon<$ 1, and a graph $G$ with $n$ vertices such that every component is either a cycle or a path. Let $c$ be the number of the cycles. The goal is to remove some of the edges such that
@@
Any LOCAL algorithm with shared randomness that solves Problem 1 with high probability, takes at least $\Omega\left(\frac{1}{\varepsilon}\right)$ rounds to finish. [^4] The lower bound is still true even if the second requirement of Problem 1 holds in expectation.
@@
We make some definitions. For a vertex $u$ in $G$, the $r$-neighborhood of $u$, denoted by $N_{r}(u)$, is the subgraph consisting of the vertices with distance at most $r$ from $u$ (identified by their vertex IDs), and the edges with at least one endpoint with distance at most $r-1$. Observe this is slightly different than the definition of $r$-neighborhood in Section 5.
@@
Take this probability to be $p$. If $p\geq 2\varepsilon$, then the algorithm does not satisfy the second requirement of Problem 1, that is does not remove at most $2+\varepsilon n$ edges with high probability. If $p<2\varepsilon$, then the algorithm does not satisfy the first requirement, since the smaller cycle has length $\frac{1}{4\varepsilon}$ and removing the edges with probability $2\varepsilon$ would mean that no edges are removed with probability $1-2\varepsilon\cdot\frac{1}{4\varepsilon}=\frac{1}{2}$. Observe, that if the algorithm solved the problem on any graph with high probability, it would have done so also on this distribution. Therefore, any algorithm that runs in $\frac{1}{10\varepsilon}$ rounds does not solve Problem 1 with high probability. ∎
@@ A.2 MPC Lower Bound
[…]
@@
Here, we review the definitions and the main result of [CzumajDP21a]. For our application, we have removed the dependence on the maximum degree from the statements.
@@
For the definition of component-stability to make sense, for each vertex, a name and an ID is defined. The names are unique but the IDs are required to be unique only within each component. The output of a component-stable algorithm may depend only on the IDs, while the names are used to distinguish between the vertices in the output.
The other definitions are rather technical and the motivations behind them lie within the proof of Theorem 4 (the main theorem of [CzumajDP21a]).
@@
Let $\delta\in(0,1)$ and $\varepsilon=\Omega\left(\frac{1}{\textnormal{poly}\log n}\right)$. Under the 1vs2-Cycle conjecture, any component-stable MPC algorithm with $O(n^{\delta})$ space per machine and total space $\textnormal{poly}(n)$ that solves the $\varepsilon$-almost-spanning forest problem (Problem 1) with probability $1-\frac{1}{n}$ must take $\Omega\left(\log\frac{1}{\varepsilon}\right)$ rounds to complete.
@@
We give a reduction from the $\varepsilon$-almost-spanning forest problem (Problem 1) to the $(1+\varepsilon)$-approximation of MST in $(1,2)$-metrics. Let $\mathcal{A}$ be an MPC algorithm as described in the statement that computes a $(1+\varepsilon/2)$-approximate MST. We derive an algorithm $\mathcal{A}^{\prime}$ for Problem 1. Take an instance $G$ for Problem 1. Construct a $(1,2)$-metric by letting all the pairs in $E(G)$ have weight $1$, and letting the rest of the pairs have weight $2$. Let $T$ be the output of $\mathcal{A}$ on the metric. We let the output of $\mathcal{A}^{\prime}$ be $F$, the set of weight-$1$ edges in $T$, i.e. $F=T\cap G$. We claim that if $T$ is a $(1+\varepsilon/2)$-approximate MST, then $F$ is an $\varepsilon$-almost-spanning forest.
First, observe that if $T$ is a $(1+\varepsilon/2)$-approximate MST, the first requirement of Problem 1 is satisfied because $F$ is a subset of a tree, therefore, a forest.
@@
Here, (8) follows from $c\leq\lvert E(G)\rvert$. This implies that $F$, in the notion of Problem 1, removes at most $c+\varepsilon(n-1)$ weight-$1$ edges, i.e. $\lvert G\setminus F\rvert\leq c+\varepsilon(n-1)$, and the second requirement is satisfied.
In conclusion, this reduction provides an algorithm $\mathcal{A}^{\prime}$ for Problem 1. Because whenever algorithm $\mathcal{A}$ outputs a $(1+\varepsilon/2)$-approximate $T$, then $\mathcal{A}^{\prime}$ outputs an $\varepsilon$-almost-spanning forest. Also, note that since $\mathcal{A}$ is component stable on the weight-$1$ edges, so is $\mathcal{A}^{\prime}$. This completes the reduction, and the lower bound is implied from Theorem 5. ∎
The remainder of this subsection is devoted to the proof of Theorem 5. We show that Theorem 4 applies to Problem 1 and Theorem 3 through a series of claims.
@@
Problem 1 can be expressed as a labeling problem.
@@
    the two requirements of Problem 1 are satisfied. ∎
@@
Problem 1 is defined on a normal family of graphs.
@@
Let $G\in\mathcal{H}$ be a graph with $n$ vertices an $c$ cycles. Let $\Gamma_{G}$ be a graph with $k\geq n$ copies of $G$ and at most $n$ extra isolated vertices. Let $L$ be valid labeling of $\Gamma_{G}$ such that every copy of $G$ has the same labeling, and let $L^{\prime}$ be the restriction of $L$ to a copy of $G$. To prove the claim, we need to show that $L^{\prime}$ is valid for $G$. To do so, we check the two requirements of Problem 1.
@@
Hence, the second requirement holds but with a parameter that is a constant factor larger, namely $2\varepsilon$. This is sufficient for the purposes of Theorem 4 (see the proof of Lemma 25 in [CzumajDP21a] this property is utilized, or Lemma 11 in the same paper for another example where the parameter is larger by a constant factor), and the problem is $1$-replicable. ∎
Putting these together gives Theorem 5.
Proof of Theorem 5.
First, we prove the theorem for $\varepsilon=\Omega\left(\frac{1}{\log^{\gamma}n}\right)$, where $\gamma\in(0,1)$. Note that $\frac{1}{\varepsilon}$ is constrained. Due to Theorem 3, we have a $\Omega\left(\frac{1}{\varepsilon}\right)$ round lower bound for Problem 1 in the randomized LOCAL model with shared randomness. Therefore, by Claims A.6, A.7 and A.8, we can apply Theorem 4 and get a $\Omega\left(\log\frac{1}{\varepsilon}\right)$ round lower bound for component-stable MPC algorithms, which concludes the proof.

### v2

@@ 5 A Lower Bound for Metric MST
Claim˜5.3 follows easily from Claim˜5.2.
Proof of Claim˜5.3.
We are given a graph $G$, comprising $\frac{1}{\varepsilon^{k}}$ cycles $C_{0},\ldots,C_{1/\varepsilon^{k}-1}$ of length $\varepsilon^{k}n$, and one vertex $u^{i}_{0}$ per cycle. Using Claim˜5.2, we compute for each vertex $u$, its $\frac{1}{\varepsilon}$-neighborhood. This neighborhood can be collected in one machine, i.e. the subpath of length $\frac{2}{\varepsilon}$ centered at the vertex is retrieved.
To create $G^{\prime}$, on the same vertex set, we connect each vertex to the two vertices at distance $\frac{1}{\varepsilon}$ from it in $G$. Since the length of the cycles is divisible by $\frac{1}{\varepsilon}$, each cycle $C_{i}$ is broken down into $\frac{1}{\varepsilon}$ cycles $C^{i}_{0},\ldots,C^{i}_{1/\varepsilon-1}$ of length $\varepsilon^{k+1}n$. Also, $u^{i}_{j}$ is already computed for $0\leq i\leq 1/\varepsilon^{k}-1$ and $0\leq j\leq 1/\varepsilon-1$ using Claim˜5.2. It is the subpath starting at $u^{i}_{0}$. ∎
Acknowledgments
Part of this work was conducted while the first and second named authors were visiting the Simons Institute for the Theory of Computing as participants in the Sublinear Algorithms program.
Appendix A Lower Bounds (LOCAL and Component-Stable MPC)
In this section, we present lower bounds for MST approximation on $(1,2)$-metrics. First, we prove a lower bound for a slightly easier problem in the LOCAL model with shared randomness. Then, we use the framework of [GKU19] and [CDP21] to lift it and get a $\Omega\left(\log\frac{1}{\varepsilon}\right)$-round lower bound for component-stable MPC algorithms that compute a $(1+\varepsilon)$-approximate MST in $(1,2)$-metrics.
We note that while the lower bound proved in this section is weaker than the lower bound of Theorem˜2 which does not require the component stability assumption, we state it nonetheless to contrast lower bounds obtained from this framework with our lower bound of Theorem˜2.
@@ A.1 LOCAL Model
We restate some of the claims and definitions from Section˜5. Given a graph $G$, the $(1,2)$-metric obtained from $G$ is such that for any two vertices $u$ and $v$, their distance in the metric is $1$ if and only if there is an edge between them in $G$. We say a spanning tree $T$ of the metric removes an edge $e\in G$ if $e\notin T$ (Definition˜5.6). That is, $T$ is constructed by removing some edges from $G$, turning it into a forest, and then connecting the resulting components with weight-$2$ edges. When $G$ is a disjoint union of cycles, a $(1+\varepsilon)$-approximate MST of this metric removes at least one edge from each cycle, and at most $c+\varepsilon(n+c-2)$ edges in total, where $n$ is the number of vertices, and $c$ is the number of cycles (Claim˜5.7).
@@
In the $\varepsilon$-almost-spanning forest problem, we are given a parameter $0<\varepsilon<1$, and a graph $G$ with $n$ vertices such that every component is either a cycle or a path. Let $c$ be the number of the cycles. The goal is to remove some of the edges such that
@@
Any LOCAL algorithm with shared randomness that solves Problem˜1 with high probability, takes at least $\Omega\left(\frac{1}{\varepsilon}\right)$ rounds to finish. [^4] The lower bound is still true even if the second requirement of Problem 1 holds in expectation.
@@
We make some definitions. For a vertex $u$ in $G$, the $r$-neighborhood of $u$, denoted by $N_{r}(u)$, is the subgraph consisting of the vertices with distance at most $r$ from $u$ (identified by their vertex IDs), and the edges with at least one endpoint with distance at most r1r-1. Observe this is slightly different than the definition of $r$-neighborhood in Section˜5.
@@
Take this probability to be $p$. If $p\geq 2\varepsilon$, then the algorithm does not satisfy the second requirement of Problem˜1, that is does not remove at most $2+\varepsilon n$ edges with high probability. If $p<2\varepsilon$, then the algorithm does not satisfy the first requirement, since the smaller cycle has length $\frac{1}{4\varepsilon}$ and removing the edges with probability $2\varepsilon$ would mean that no edges are removed with probability $1-2\varepsilon\cdot\frac{1}{4\varepsilon}=\frac{1}{2}$. Observe, that if the algorithm solved the problem on any graph with high probability, it would have done so also on this distribution. Therefore, any algorithm that runs in $\frac{1}{10\varepsilon}$ rounds does not solve Problem˜1 with high probability. ∎
@@ A.2 MPC Lower Bound
In this subsection, we prove a lower bound for Problem˜1 in the MPC model by using the framework of [GKU19] and [CDP21] to lift the lower bound in Section˜A.1.
@@
Here, we review the definitions and the main result of [CDP21]. For our application, we have removed the dependence on the maximum degree from the statements.
@@
For the definition of component-stability to make sense, for each vertex, a name and an ID is defined. The names are unique, but the IDs are required to be unique only within each component. The output of a component-stable algorithm may depend only on the IDs, while the names are used to distinguish between the vertices in the output.
The other definitions are rather technical, and the motivations behind them lie within the proof of Theorem˜4 (the main theorem of [CDP21]).
@@
Let $\delta\in(0,1)$ and $\varepsilon=\Omega\left(\frac{1}{\textnormal{poly}\log n}\right)$. Under the 1vs2-Cycle conjecture, any component-stable MPC algorithm with $O(n^{\delta})$ space per machine and total space $\textnormal{poly}(n)$ that solves the $\varepsilon$-almost-spanning forest problem (Problem˜1) with probability $1-\frac{1}{n}$ must take $\Omega\left(\log\frac{1}{\varepsilon}\right)$ rounds to complete.
@@
We give a reduction from the $\varepsilon$-almost-spanning forest problem (Problem˜1) to the $(1+\varepsilon)$-approximation of MST in $(1,2)$-metrics. Let $\mathcal{A}$ be an MPC algorithm as described in the statement that computes a $(1+\varepsilon/2)$-approximate MST. We derive an algorithm $\mathcal{A}^{\prime}$ for Problem˜1. Take an instance $G$ for Problem˜1. Construct a $(1,2)$-metric by letting all the pairs in $E(G)$ have weight $1$, and letting the rest of the pairs have weight $2$. Let $T$ be the output of $\mathcal{A}$ on the metric. We let the output of $\mathcal{A}^{\prime}$ be $F$, the set of weight-$1$ edges in $T$, i.e. $F=T\cap G$. We claim that if $T$ is a $(1+\varepsilon/2)$-approximate MST, then $F$ is an $\varepsilon$-almost-spanning forest.
First, observe that if $T$ is a $(1+\varepsilon/2)$-approximate MST, the first requirement of Problem˜1 is satisfied because $F$ is a subset of a tree, therefore, a forest.
@@
Here, (8) follows from $c\leq\lvert E(G)\rvert$. This implies that $F$, in the notion of Problem˜1, removes at most $c+\varepsilon(n-1)$ weight-11 edges, i.e. $\lvert G\setminus F\rvert\leq c+\varepsilon(n-1)$, and the second requirement is satisfied.
In conclusion, this reduction provides an algorithm $\mathcal{A}^{\prime}$ for Problem˜1. Because whenever algorithm $\mathcal{A}$ outputs a $(1+\varepsilon/2)$-approximate $T$, then $\mathcal{A}^{\prime}$ outputs an $\varepsilon$-almost-spanning forest. Also, note that since $\mathcal{A}$ is component-stable on the weight-$1$ edges, so is $\mathcal{A}^{\prime}$. This completes the reduction, and the lower bound is implied from Theorem˜5. ∎
The remainder of this subsection is devoted to the proof of Theorem˜5. We show that Theorem˜4 applies to Problem˜1 and Theorem˜3 through a series of claims.
@@
Problem˜1 can be expressed as a labeling problem.
@@
    the two requirements of Problem˜1 are satisfied. ∎
@@
Problem˜1 is defined on a normal family of graphs.
@@
Let $G\in\mathcal{H}$ be a graph with $n$ vertices an $c$ cycles. Let $\Gamma_{G}$ be a graph with $k\geq n$ copies of $G$ and at most $n$ extra isolated vertices. Let $L$ be valid labeling of $\Gamma_{G}$ such that every copy of $G$ has the same labeling, and let $L^{\prime}$ be the restriction of $L$ to a copy of $G$. To prove the claim, we need to show that $L^{\prime}$ is valid for $G$. To do so, we check the two requirements of Problem˜1.
@@
Hence, the second requirement holds but with a parameter that is a constant factor larger, namely $2\varepsilon$. This is sufficient for the purposes of Theorem˜4 (see the proof of Lemma 25 in [CDP21] where this property is utilized, or Lemma 11 in the same paper for another example where the parameter is larger by a constant factor), and the problem is $1$-replicable. ∎
Putting these together gives Theorem˜5.
Proof of Theorem˜5.
First, we prove the theorem for $\varepsilon=\Omega\left(\frac{1}{\log^{\gamma}n}\right)$, where $\gamma\in(0,1)$. Note that $\frac{1}{\varepsilon}$ is constrained. Due to Theorem˜3, we have a $\Omega\left(\frac{1}{\varepsilon}\right)$ round lower bound for Problem˜1 in the randomized LOCAL model with shared randomness. Therefore, by Claims˜A.6, A.7 and A.8, we can apply Theorem˜4 and get a $\Omega\left(\log\frac{1}{\varepsilon}\right)$ round lower bound for component-stable MPC algorithms, which concludes the proof.
@@
Appendix B Implications for TSP
In the traveling salesman problem (TSP), given a weighted graph $G$, the goal is to find the shortest cycle that includes all the vertices (exactly once). When combined with the techniques of [JMNZ24], Theorem˜1 leads to the following corollary for TSP.
Theorem 7.
Given a metric, for any fixed $\delta\in(0,1)$ and any $\varepsilon>0$, a $(2+\varepsilon)$-approximate TSP can be computed in $O\left(\log\frac{1}{\varepsilon}+\log\log n\right)$ rounds of the MPC model, with $O(n^{\delta})$ space per machine and $\widetilde{\Theta}(n^{2})$ total space.
Proof.
In metrics, any closed walk can be converted to a closed cycle with the same vertex set and at most the same weight. To do so, it suffices to go over the vertices of the walk one by one and append each vertex to the output sequence if it has not already been added. As the distances between the vertices satisfy the triangle inequality, skipping over the repeated vertices does not increase the cost. Therefore, the weight of the resulting cycle is not larger than the weight of the original walk. This can be implemented in the MPC model in $O(1)$ rounds.
Since $\textnormal{MST}\leq\textnormal{TSP}$, any Eulerian tour of a $(1+\varepsilon)$-approximate MST provides a $(2+2\varepsilon)$-approximate TSP, where an Eulerian tour of a tree $T$ is a closed walk on the edges of $T$ that includes each edge exactly twice. [JMNZ24] showed that the Eulerian tour of a tree can be computed in few rounds of the MPC model provided that the tree satisfies certain hierarchical properties:
Lemma B.1 (Theorem 16 of [JMNZ24]).
Given a set of vertices $V,$ let $\mathcal{P}_{0},\mathcal{P}_{1},\ldots,\mathcal{P}_{L}$ be a hierarchy of partitions such that $\mathcal{P}_{0}=\{\{u\}\mid u\in V\}$ and $\mathcal{P}_{L}=\{V\}$. Let $T=\bigcup_{1\leq i\leq L}E_{i}$ be a tree such that $\mathcal{P}_{i}=\mathcal{P}_{i-1}\oplus E_{i}$ and $\lvert\mathcal{P}_{i-1}\rvert-\lvert\mathcal{P}_{i}\rvert=\lvert E_{i}\rvert$. For each level $i$, consider the forest $F_{i}$ obtained by taking $E_{i}$ and contracting all the vertex sets of $\mathcal{P}_{i-1}$, and let $\Lambda$ be an upperbound for the unweighted diameter of the components in $F_{i}$. There exists an MPC algorithm that computes an Eulerian tour of $T$ in $O(\log L+\log\Lambda)$ rounds, using total space $O(nL+n^{1+\delta})$ and space $O(n^{\delta})$ per machine, where $\delta\in(0,1)$ is an arbitrary constant.
The hierarchy $\{\widehat{\mathcal{P}}_{t}\}_{t=\alpha^{k}}$ and the edge sets $\{E_{t}\}_{t=\alpha^{k}}$, computed by Algorithm˜1, satisfy the hierarchical requirements. The number of levels is $L=\log_{\alpha}n=O(\log n)$. To bound $\Lambda$, observe that for each level $t=\alpha^{k}$, $E_{t}$ is obtained by performing $O\left(\log\frac{1}{\varepsilon}+\log\log n\right)$ rounds of the modified Borůvka algorithm where each round consists of computing and contracting a set of stars (step 1), and the edges computed in step 1 can be taken to be a star. Therefore, $\Lambda=\exp\left(O\left(\log\frac{1}{\varepsilon}+\log\log n\right)\right)$.
To conclude the proof, we can first compute the $(1+\varepsilon)$-approximate MST along with a hierarchy of partitions by utilizing Algorithm˜1, in $O\left(\log\frac{1}{\varepsilon}+\log\log n\right)$ rounds, using total space $\widetilde{O}(n^{2})$. Then, by applying Lemma˜B.1, we can obtain the $(2+\varepsilon)$-approximate TSP, in $O\left(\log\frac{1}{\varepsilon}+\log\log n\right)$ rounds, using total space $O(n^{1+\delta})$. That is $O\left(\log\frac{1}{\varepsilon}+\log\log n\right)$ rounds and $\widetilde{O}(n^{2})$ total space overall, as desired. ∎